OpenStudy (anonymous):

Which answer shows this equation in standard form? 8 – 4(x – y) = –6x + 6

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

First distribute -4 into the parenthesis.

OpenStudy (anonymous):

8-4(-4-y)=-6x+6

OpenStudy (igreen):

\(-4(x - y)\) \(-4 \times x + -4 \times y\) \(-4x + 4y\) So we have: \(8 - 4x + 4y = -6x + 6\) Subtract 8 to both sides: \(-4x + 4y = -6x - 2\) Now add 6x to both sides, what do you get?

OpenStudy (anonymous):

2+y=-6x-14

OpenStudy (igreen):

Um..no..

OpenStudy (igreen):

Add \(-4x + 6x\) tell me what you get.

OpenStudy (anonymous):

2x

OpenStudy (igreen):

Yes, so that gives us: \(2x + 4y = -2\)

OpenStudy (anonymous):

Thank you!